OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 4) with a slope of –1 in standard form. –x + 4y = –1 x – 4y = –1 x + y = 3 x – y = –3

OpenStudy (anonymous):

@Prateek_agr Your The Best! : )

OpenStudy (anonymous):

y=-x+b x=-1 then y =4: 4 = 1 + b => b=3 So, y=-x + 3 or x + y = 3 is the answer.

OpenStudy (anonymous):

Thank you friend :)